The length of one leg of a right triangle is 3x the length of the other. The length of the hypotenuse is 10. What is the length of the longest leg?
Answer by jorel1380(3719)

If one leg of a right triangle is three times the other, let X be the length of the short leg. Then the long leg is 3X. In this case:
a^2+b^2=c^2
X^2+(3X)^2 = 10^2
X^2+9X^2 = 100
10X^2 = 100
X^2 = 10
X = sq.rt.(10)
3X = 3 sq.rt.(10)